Find a17 of arithmetic sequence (2,-24,-50,-76)
1
Expert's answer
2021-10-12T02:11:44-0400

Arithmetic sequence nth theorem formula

"a_n=a_1+(n-1)d\\
a_6=24=a_1+(6-1)d \implies a_1+5d=24......(1)\\
a_{11}=39=a_{11}+(11-1)d \implies a_1+10d=39......(2)\\
Adding \space 1 \& 2\\
d= 3\\
a_1=9\\
So, a_{17}=9+(17-1)*3=57"